Please help me solve this equation: If angle A is 8 degrees less than three times angle B. If side by side they form a straight angle, find each angle.
Answer by macston(5194)

A=3B-8
A+B=180 degrees
Substitute for A:
3B-8+B=180
3B+B=188
4B=188
B=47 ANSWER 1
A=3(47)-8
A=133 ANSWER 2
Check:
A+B=180
133+47=180
180=180